# QRazor: Reliable and Effortless 4-bit LLM Quantization by Significant Data Razoring

### v1

@@ A.2 Importance of Weight Quantization
A.3 SDR Encoding: Detecting the Razoring Point with Bitwise OR Operations
@@
A.4 Perplexity
We evaluated the performance of our QRazor scheme using perplexity metrics for W4A4, W4A8, W4A4KV4, and W4A8KV4 configurations on the LLaMA-2-7B and LLaMA-2-13B models, specifically on the Lambada-OpenAI dataset (Paperno et al., 2016). For calibration, 128 randomly selected samples from wikitext2 data has been used.
@@ A.4 Perplexity
A.5 Overall Accuracy Result
@@
A.6 FLOPS & OPS Comparison Result
To evaluate the computational efficiency of our proposed quantization scheme relative to the approach outlined in the QuaRot paper, we conducted a comparative analysis of the FLOPs and OPs involved in the attention layer of a transformer architecture. QuaRot employs two primary operations—Hadamard product and rotation operation—for outlier-aware quantization. These operations effectively reduce the impact of outliers during quantization by transforming datasets into lower-variance distributions. This process enables efficient mapping of data to 4-bit representations using conventional per-tensor or per-channel quantization techniques.
However, the Hadamard product and rotation operations introduce substantial computational overhead during both quantization and dequantization. Notably, the dequantization process in QuaRot incurs a significant number of additional FLOPs, as it involves dequantization computations for weights as well. Consequently, while QuaRot is effective for quantization, its reliance on FLOPs-intensive operations reduces computational efficiency when compared to other methods with similar OPs.
In contrast, our proposed quantization scheme simplifies the dequantization process by utilizing shifting operations immediately after INT matrix-multiplication(matmul) to perform decompression. This design choice results in a 35% increase in OPs relative to QuaRot. However, it dramatically reduces the FLOPs required, demanding only 3% of the FLOPs utilized by QuaRot. This substantial reduction in FLOPs contributes to increased throughput and lower power consumption, underscoring the computational efficiency and practical benefits of our method compared to QuaRot.
Figure 5: FLOPs & OPs comparsion with QuaRot in Attention Layer
A.7 Accuracy Comparison with SOTA work
@@
A.8 QRazor Quantization Flow in Transformers
@@
Figure 6: QRazor Quantiztaion Flow in Attention & FFN Layer

### v2

@@ A.1 Importance of the Weight Quantization
A.2 SDR Encoding: Detecting the Razoring Point with Bitwise OR Operations
@@
A.3 Perplexity
We evaluated the performance of our QRazor scheme using additional perplexity metric of Lambada-OpenAI dataset (Paperno et al., 2016) for W4A4, W4A8, W4A4KV4, and W4A8KV4 configurations on the LLaMA-2-7B and LLaMA-2-13B models. For calibration, 128 randomly selected samples from wikitext2 data has been used.
@@ A.3 Perplexity
A.4 FLOPs & OPs Comparison Results
To evaluate the computational efficiency of our proposed quantization scheme relative to the approach outlined in the QuaRot paper, we conducted a comparative analysis of the FLOPs and IOPs involved in the attention layer of a transformer architecture. QuaRot employs two primary operations—Hadamard product and rotation operation—for outlier-aware quantization. These operations effectively reduce the impact of outliers during quantization by transforming datasets into lower-variance distributions. This process enables efficient mapping of data to 4-bit representations using conventional per-tensor or per-channel quantization techniques.
However, the Hadamard product and rotation operations introduce substantial computational overhead during both quantization and dequantization. Notably, the dequantization process in QuaRot incurs a significant number of additional FLOPs, as it involves dequantization computations for weights as well. Consequently, while QuaRot is effective for quantization, its reliance on FLOPs-intensive operations reduces computational efficiency when compared to other methods with similar IOPs.
In contrast, our proposed quantization scheme simplifies the dequantization process by utilizing shifting operations immediately after INT matrix-multiplication(matmul) to perform decompression.
Table 8: OPs comparison table of rotation and SDR computations
Metric Hadamard (Rotation) QRazor (SDR)
Operation Single Hadamard Hadamard Heads SDR Compression Barrel Shifter
Formula $M\times N$ $H\times M\times N$ ($M\times N\times 2$)/$G$ ($M\times N$)/$G$
Result ($M=128,N=64,H=8,G=32$) $8,192$ FLOPs $65,536$ FLOPs $512$ IOPs $256$ IOPs
A.5 Overall Accuracy Results
@@
A.6 Accuracy Comparison with SOTA Works
@@
A.7 The Quantization Flow of QRazor in Transformers
@@
Figure 5: The QRazor quantization flow in attention and feed-forward network layers